β-Galactosidase (β-gal) is a hydrolase enzyme that catalyzes the hydrolysis of β-galactosides into monosaccharides. A 0.360 gram sample of β-galactosidase is dissolved in water to make 0.108 L of solution, and the osmotic pressure of the solution at 25 °C is found to be 0.709 mbar. Calculate the molecular mass of β-galactosidase.
Answer #1

The concentration of the solution is calculated from the osmotic pressure equation:

C = Po / R * T = (0.709 bar * 1 atm / 1.013 bar) / 0.082 atm * L / mol * K * (25 + 273) K = 0.0286 M

The moles are calculated:

n = M * V = 0.0286 M * 0.108 L = 3.09x10 ^ -3 mol

The molar mass is calculated:

MM = g / n = 0.360 g / 3.09x10 ^ -3 mol = 116.5 g / mol

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.